OpenStudy (anonymous):

Solve for the variable: 2/3 x =26

OpenStudy (anonymous):

2/3x = 26 divide by 2/3 divide by 2/3

OpenStudy (anonymous):

=38.81

OpenStudy (anonymous):

No. Multiply by the reciprocal 3/2 so (3/2) (2/3)x= 26 (3/2) so then 26*3/2= 39Thus x=39

OpenStudy (anonymous):

is it (2/3)x = 26 or 2/(3x) = 26 ?

OpenStudy (anonymous):

My answer is assuming (2/3)x = 26

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x =26\] \[(=) 2x=26*3 (=) x=\frac{ 26*3 }{ 2 } (=) x=39\] this is how i would do it but the answer is still 39